# rtboxstyle linecolor=black, backgroundcolor=lightlight-gray, leftmargin=0cm, rightmargin=0cm, skipabove=4pt, skipbelow=4pt, linewidth=0.4pt, innertopmargin=3pt, innerbottommargin=3pt, innerrightmargin=6pt, innerleftmargin=6pt, roundcorner=2

### v1

@@ 5.3 Semantic Correctness Evaluation
We use the semantic correctness validation mechanism introduced in Section 4.6 to evaluate the semantic of synthetic graphs. In the experiment, the GAT model was trained for 2500 epochs with a learning rate of 0.0001, while the MLP classifier was trained for 1500 epochs with the same learning rate. To validate the effectiveness and interpretability of our semantic accuracy evaluation, we tested it on real-world provenance graphs from the same datasets but different from the training graphs. The results show that the semantic accuracy of these real graphs exceeds 90% across the Nodlink, Cadets, Theia, and Trace datasets. This supports the validity of our semantic accuracy evaluation method. The results in Figure 2 show that the provenance graphs generated by ProvSyn consistently outperform the baselines in semantic accuracy across all four datasets, showing robustness to differences in data distribution. In contrast, the baselines exhibit substantial fluctuations in semantic accuracy across datasets. Notably, in the more challenging Theia and Trace datasets, the semantic accuracy of baseline-generated graphs remained below 0.20, whereas ProvSyn achieved 0.57 and 0.82, respectively. These results demonstrate the robustness of ProvSyn in handling complex, low signal-to-noise provenance data.
5.4 Data Imbalance Evaluation
To evaluate whether the inclusion of ProvSyn-generated synthetic graphs mitigates class imbalance, we use two metrics—entropy and the Gini index—to quantify the distributional balance of node and edge labels. Given a label distribution with probabilities $p_{1},p_{2},\ldots,p_{n}$, the two metrics are defined as:
$\text{Entropy}=-\sum_{i=1}^{n}p_{i}\log(p_{i}),\quad\text{Gini}=1-\sum_{i=1}^{n}p_{i}^{2}$
Higher values of both metrics indicate a more uniform and balanced label distribution.
We conduct experiments on the Nodlink, Cadets, Theia, and Trace datasets, both in their original data and after incorporating ProvSyn-generated data. For each dataset, we add 100 synthetic graphs merged with the original graph as graph communities. The results are summarized in Table V. As shown, the inclusion of ProvSyn increases both entropy and Gini index for node and edge labels across all datasets. In the Nodlink dataset, the metrics improve significantly due to the dataset’s small scale. These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes in provenance graph datasets.
To evaluate whether the synthesized graphs extend beyond the training data distribution and contain novel structures that cover more provenance scenarios, we use two metrics: Novelty and Uniqueness. Novelty is computed by verifying that each generated graph is neither a subgraph of any reference graph in the training set nor contains any reference graph as a subgraph, indicating whether novel structures exist in the synthesized graphs. Uniqueness measures the proportion of generated graphs that are structurally distinct from one another, reflecting the diversity of the generated graphs. As shown in Table VI, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets. This indicates that the synthetic data exhibits strong diversity and effectively captures patterns that were underrepresented or missing in the original data.
TABLE V: Comparison of node and edge label entropy and Gini index across different datasets with and without synthetic data. Higher values indicate better performance.
Dataset Data Source Node Label Edge Label
Entropy Gini Index Entropy Gini Index
Nodlink Real 1.03 0.42 2.41 0.75
+ProvSyn 1.31 (+0.28) 0.55 (+0.13) 2.63 (+0.22) 0.82 (+0.07)
Cadets Real 1.11 0.44 1.84 0.52
+ProvSyn 1.31 (+0.20) 0.53 (+0.09) 1.92 (+0.08) 0.54 (+0.02)
Theia Real 1.31 0.53 1.33 0.39
+ProvSyn 1.41 (+0.10) 0.58 (+0.05) 1.36 (+0.03) 0.39 (+0.00)
Trace Real 1.52 0.57 1.57 0.45
+ProvSyn 1.62 (+0.10) 0.61 (+0.04) 1.61 (+0.04) 0.47 (+0.02)
TABLE VI: Evaluation of ProvSyn-generated graphs in terms of Novelty and Uniqueness across different datasets. Higher values indicate better performance.
Dataset Novelty (%) Uniqueness (%)
Nodlink 100.0 100.0
Cadets 100.0 100.0
Theia 100.0 100.0
Trace 100.0 100.0
@@ 5.5 Application in IDS
To demonstrate the practical utility of ProvSyn, we focus on the task of APT attack detection. We conduct experiments using two detection algorithms, Magic[31] and Nodlink[37]. Specifically, we augment the training set with synthesized provenance graphs and then observe whether the detection performance improves or declines during testing.
For the Nodlink algorithm, we enhanced its dataset by adding 50 synthetic graphs merged with the original provenance graph as graph communities. This resulted in a 15% increase in node counts. For the Magic algorithm, we enhanced the Cadets, Theia, and Trace datasets. Due to their larger scale, we added 300 synthetic graphs merged into their original provenance graphs. This led to a 10% to 15% increase in node counts.
To evaluate the effectiveness of synthetic graphs in intrusion detection tasks, we maintain identical training configurations before and after incorporating synthetic data. The detection performance is presented in Table VII. In addition to Precision and Recall, we report the number of False Positives (#FP), an crucial metric for intrusion detection system, as a high false positive rate can increase the workload of subsequent security analysis. As shown in the results, on the Nodlink dataset, the inclusion of synthetic graphs leads to no change in Precision, Recall, or the number of false positives, indicating that the synthetic data is of comparable quality to the original data. On the Cadets, Theia, and Trace datasets, Recall remains constant, and Precision improves. Specifically, false positives decrease by 506 on Cadets (from 6435 to 5929), by 95 on Theia (from 931 to 836), and by 379 on Trace (from 2390 to 2011). These results suggest that augmenting with synthetic graphs produced by ProvSyn improves model performance, and effectively reduces false positives in detection systems.
TABLE VII: IDS performance across different datasets with and without synthetic data. Higher values indicate better performance.
Dataset Data Source Precision(%) Recall(%) #FP
Nodlink Real 100.00 100.00 0
+ ProvSyn 100.00 100.00 0
Cadets Real 66.58 99.79 6435
+ ProvSyn 68.37 99.77 5929
Trace Real 96.61 99.98 2390
+ ProvSyn 97.13 99.98 2011
Theia Real 96.45 99.99 931
+ ProvSyn 96.80 99.99 836
@@ 5.6 Overhead Analysis
TABLE VIII: Time and memory requirements of ProvSyn’s core modules.
Module Phase Time Memory
Graph Generation Training 18 hours 0.84 GB
Inference 28.09 s 0.38 GB
Text Generation Training 109 s 3.44 GB
Inference 2.78 hours 28 GB
In this section, we analyze the time and memory requirements of ProvSyn (see Table VIII). The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph generation module and the text attribute generation module.
For the heterogeneous graph generation module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs. The average training time is 18 hours, with a GPU memory usage of 0.84 GB. Inference for generating 1,000 graphs takes 28.09 seconds and consumes 0.38 GB of memory.
For the text attribute generation module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization. The FastLanguageModel library is used to accelerate training. The training consists of 60 steps, takes an average of 109 seconds, and requires a peak memory of 3.44 GB. During inference, generating text attributes for a single graph takes an average of 80 seconds. In our experiments, we deploy 8 fine-tuned LLMs in parallel to synthesize 1,000 graphs, completing the process in 2.78 hours. The total GPU memory required for the 8 models is 28 GB.
This analysis demonstrates that ProvSyn is a time- and memory-efficient framework for provenance graph synthesis.
@@ 6 Discussion and Limitations
Temporal Modeling. A provenance graph is not only a directed heterogeneous graph with textual attributes, but also a temporal graph in which each event is associated with a timestamp indicating its occurrence time in the system. These temporal information has recently been leveraged by downstream intrusion detection systems [53] to better capture the patterns of benign and anomalous behaviors. However, in ProvSyn, we do not synthesize explicit timestamps for events. Instead, we approximate temporal order by relying on the sequence of directed edges. Future work may explore modeling temporal dynamics using temporal graph generation networks [75], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
Graph Scale. In ProvSyn, the generated provenance graphs are smaller in scale compared to real-world provenance graphs. To enable data augmentation, we embed multiple synthetic graphs as distinct communities within large real-world provenance graphs. The results show that this approach effectively mitigates data imbalance and reduces the false positive rate of detection models trained on the augmented datasets. Future work can focus on scaling up synthetic provenance graphs by enhancing graph generation models for large-scale synthesis and exploring graph merging algorithms that combine smaller graphs into larger ones while preserving structural fidelity.
Efficiency of Text Attribute Synthesis. During textual attribute synthesis, overlaps among nodes in different sequences generated by DFS can lead to naming conflicts when these sequences are processed in parallel to generate node names. To address this issue, we serialize the sequences and input them into the LLM sequentially. The generated node names from earlier sequences are then used as context for subsequent sequences. This approach is effective for small-scale graphs. However, the synthesis efficiency decreases for large-scale graphs. Future work may explore strategies to partition large graphs into batches of non-overlapping sequences, enabling parallel processing while avoiding node naming conflicts.
@@ 7 Conclusion
This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three stages: heterogeneous graph structure generation, rule-based topological refinement, and LLM-based synthesis of textual attributes. To evaluate the fidelity of the generated graphs, we assess structural, textual, temporal, embedding, and semantic aspects. Compared with provenance data directly synthesized by LLMs, ProvSyn consistently produces graphs with higher fidelity. When integrated into existing datasets, ProvSyn-generated graphs help mitigate class imbalance and lead to reduced false positive rates in detection models trained on the augmented data.
Appendix A Entity and Event Distribution in Our Datasets
@@
We present the distributions of entity type and event type across four commonly-used public provenance datasets in Figure 3. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
Appendix B Textual Quality Evaluation
In Algorithm 2, we compute the maximum similarity score between the text attribute of each node and the reference corpus. This metric serves to evaluate the model’s maximum potential in generating accurate and relevant attributes.
Input : Generated graph nodes $V_{\text{gen}}=\{g_{1},g_{2},...,g_{m}\}$ with predicted names and types;
Real graph nodes $V_{\text{real}}=\{r_{1},r_{2},...,r_{n}\}$ with ground-truth names and types;
Node type set $\mathcal{T}=\{t_{1},t_{2},...,t_{k}\}$
Output : Final textual quality score $S_{\text{final}}$
1
2foreach $t\in\mathcal{T}$ do
       $\mathcal{R}_{t}\leftarrow\{r_{j}\mid r_{j}\in V_{\text{real}},\text{type}(r_{j})=t\}$
        // Construct reference corpus
3
[…]
6
7foreach $g_{i}\in V_{\text{gen}}$ do
8       $t\leftarrow\text{type}(g_{i})$
[…]
10       foreach $r_{j}\in\mathcal{R}_{t}$ do
[…]
14
[…]
17
[…]
  // Aggregate final score
18
19 return $S_{\text{final}}$
20
Algorithm 2 Textual Quality Evaluation
Appendix C Hyperparameters Setting
Figure 3: Entity Type and Event Type Distribution in Provenance Dataset including Nodlink, Cadets, Theia and Trace.
Figure 4: Hidden size setting in Nodlink dataset.
Figure 5: Inference temperature setting in Nodlink dataset.

### v2

@@ 5.2 Fidelity Evaluation
5.3 Data Imbalance Mitigation
Table 3: Evaluation of data imbalance mitigation. The left section includes two metrics: Novelty (Nov.) and Uniqueness (Uniq.). The middle and right sections compare Node/Edge Normalized Entropy (Real vs. Real + Synthetic), where higher values indicate better balance.
Dataset Nov. & Uniq. Node Entropy Edge Entropy
Nov. Uniq. Real +ProvSyn Real +ProvSyn
Cadets-E3 100.0 100.0 0.34 0.64 0.40 0.46
Theia-E3 100.0 100.0 0.91 0.92 0.59 0.59
ClearScope-E5 100.0 100.0 0.30 0.41 0.33 0.34
Theia-E5 100.0 100.0 0.24 0.59 0.35 0.47
OPTC-H201 100.0 100.0 0.43 0.70 0.56 0.57
OPTC-H501 100.0 100.0 0.41 0.58 0.55 0.55
Novelty and Uniqueness. As shown in Table 3, the graphs generated by ProvSyn achieve 100% novelty and 100% uniqueness across all four datasets. This indicates that the synthetic data exhibits strong diversity and effectively captures patterns that were underrepresented or missing in the original data.
Normalized Entropy. We compute normalized entropy for both the original datasets and the augmented datasets after incorporating ProvSyn-generated graphs. The results are summarized in Table 3. Overall, the inclusion of ProvSyn-generated data consistently increases the normalized entropy of node and edge label distributions across all datasets. Notably, datasets with more severe initial imbalance exhibit more pronounced entropy improvements after augmentation. For example, in Theia-E5, the node label normalized entropy increases substantially from 0.24 to 0.59. These results suggest that ProvSyn contributes to a more balanced label distribution by enriching underrepresented classes.
@@ 5.4 Application in IDS
Table 4: Performance comparison of IDS models trained on different data compositions. The “Real” rows represent models trained solely on real data, while “+ ProvSyn” rows represent models trained on a combination of real and synthesized provenance data. Green cells indicate performance improvement over the baseline, and Red cells indicate a decrease in performance. (F-Score and MCC are scaled by $10^{3}$)
Nodlink Flash
Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
Cadets-E3 Real 1.72 0.75 19.83 9.74 0.00 0.50 0.00 0
+ ProvSyn 1.94 0.84 25.41 35.24 4.85 0.59 20.47 12.55
Theia-E3 Real 0.11 0.47 -2.55 0.31 0.22 0.44 -3.24 0.56
+ ProvSyn 0.60 0.53 2.71 1.89 0.44 0.53 1.99 1.41
Clearscope-E5 Real 0.29 0.47 -3.27 0.41 0.58 0.48 -1.44 0.82
+ ProvSyn 0.97 0.61 7.93 2.50 0.83 0.57 5.53 2.00
Theia-E5 Real 1.16 0.73 15.56 12.51 0.13 0.35 -14.90 0.08
+ ProvSyn 1.03 0.80 17.20 18.63 0.30 0.60 4.08 2.29
OpTC-H201 Real 32.24 0.72 671.63 15.61 0.11 0.17 N/A 0.01
+ ProvSyn 32.61 0.72 1446.15 15.81 5.12 0.55 293.88 1.53
OpTC-H501 Real 6.10 0.70 N/A 10.73 0.15 0.26 -335.60 0.07
+ ProvSyn 6.82 0.72 225.50 12.89 0.86 0.46 -35.13 0.73
Magic Velox
Dataset Source F-Score B. Acc MCC DOR F-Score B. Acc MCC DOR
Cadets-E3 Real 0.99 0.73 14.61 12.92 0.00 0.50 -0.04 0
+ ProvSyn 1.54 0.80 20.76 17.49 3.48 0.51 6.43 7.88
Theia-E3 Real 0.42 0.57 N/A 2.00 54.61 0.57 67.78 243.04
+ ProvSyn 0.63 0.72 N/A 12.46 140.35 0.55 151.65 1934.86
Clearscope-E5 Real 0.63 0.48 -1.20 0.88 48.19 0.52 49.25 205.28
+ ProvSyn 0.86 0.61 9.32 15.30 174.76 0.59 174.48 751.84
Theia-E5 Real 0.30 0.69 7.45 44.23 54.79 0.51 120.37 11192.40
+ ProvSyn 0.39 0.76 10.07 79.14 56.34 0.51 170.24 $\infty$
OpTC-H201 Real 1.67 0.27 -1882.50 0.13 0.68 0.50 3.61 21.48
+ ProvSyn 1.67 0.28 -915.75 0.15 0.69 0.50 4.82 38.01
OpTC-H501 Real 0.42 0.26 N/A 0.11 2.65 0.50 13.77 331.65
+ ProvSyn 2.04 0.63 N/A 3.04 2.66 0.50 21.07 994.95
To demonstrate the practical utility of ProvSyn, we focus on APT attack detection using four representative systems: Nodlink [li2023nodlink], Flash [rehman2024flash], Magic [jia2024magic], and Velox [bilot2025sometimes]. All systems are reproduced using the open-source PIDSMaker [bilot2026pidsmaker, bilot2025sometimes] framework. Our evaluation focuses on whether augmenting the training set with synthesized provenance graphs enhances detection performance. To this end, we incorporate synthetic graphs, as graph communities, into every time window of the training data from the first day. The detection performance is presented in Table 4. Based on the results, we have the following four key findings.
Detection Performance Enhancement. The integration of ProvSyn data consistently strengthens detection metrics across various baselines, demonstrating its efficacy in both high-performing and low-performing scenarios. For instance, when a model already exhibits reasonable performance, such as Nodlink on the Cadets-E3 dataset, the B. Acc increases from $0.75$ to $0.84$ and the F-Score improves from $1.72\times 10^{-3}$ to $1.94\times 10^{-3}$. More importantly, the augmentation provides a critical correction in cases where the initial performance is near or below random chance. A prominent example is observed with the Flash algorithm on the OpTC-H201 dataset, where the B. Acc rises significantly from $0.17$ to $0.55$. Similar improvements are seen in the OpTC-H501 dataset for Flash, where the B. Acc progresses from $0.26$ to $0.46$, effectively rescuing the model from a state of total detection failure.
Discriminative Reliability and Diagnostic Utility. The impact of ProvSyn is further evidenced by the improvement in MCC and DOR, which reflect the model’s correlation with the ground truth and its diagnostic strength. The MCC values are consistently pulled toward positive territory, particularly in instances where the baseline models initially yield negative results, such as Nodlink on Theia-E3 moving from 2.55-2.55 to $2.71$ and Flash on Clearscope-E5 shifting from 1.44-1.44 to $5.53$. This indicates a fundamental shift from biased or inverse classification to a correct positive correlation. Simultaneously, the DOR increases across all experiments, often by several orders of magnitude. For example, the DOR for Velox on Theia-E3 jumps from $243.04$ to $1934.86$, and on Theia-E5, it reaches $\infty$. These results suggest that the diagnostic efficiency for identifying attack samples is substantially sharpened through the augmented training set.
Multi-Factor Sensitivity Analysis. The efficacy of ProvSyn is closely coupled with three key factors. Regarding the IDS architecture, the improvement for Flash on OpTC-H201 is more pronounced than for Nodlink, which may be because the baseline detection performance of Flash is lower than that of Nodlink, providing a larger margin for potential enhancement. For dataset characteristics, while Magic shows negligible F-Score improvement on OpTC-H201, it demonstrates significant F-score improvements across the other five datasets. Finally, metric-specific sensitivity is observed in Velox; while its B.Acc on OpTC-H201 remains stable at $0.50$, its MCC and DOR show notable increases from $3.61$ to $4.82$ and $21.48$ to $38.01$ respectively, suggesting that while the absolute detection threshold may be constrained by the algorithm’s architecture, its discriminative confidence and diagnostic power are still bolstered by the augmented data.
Enhancement Mechanism Analysis. The overall performance uplift across different dataset–IDS combinations can be attributed to the mitigation of data imbalance and the diversity introduced by ProvSyn synthesized samples. By incorporating a wider variety of benign provenance graphs into the training set, the models can cover previously underrepresented system behaviors and edge-case scenarios. This expanded representation of the benign space helps the detection algorithms to more accurately define the boundaries of normal activity. Consequently, the models become less prone to overfitting on limited real-world benign data, enabling them to better distinguish subtle malicious deviations from legitimate but rare system actions, thereby improving the robustness of APT detection.
@@ 5.5 Overhead Analysis
We analyze the time and memory requirements of ProvSyn. The most time- and resource-intensive components of the ProvSyn framework are the heterogeneous graph synthesis module and the text attribute synthesis module.
For the graph structure synthesis module, we use 16,000 samples for training, with a batch size of 32 and 3,000 training epochs. The average training time is 18 hours, with a GPU memory usage of 0.84 GB. Inference for generating 1,000 graphs takes 28.09 seconds and consumes 0.38 GB of memory. For the text attribute synthesis module, we adopt Llama3.2-3B-Instruct as the base model, applying the LoRA algorithm with 4-bit quantization. The FastLanguageModel library is used to accelerate training. The training consists of 60 steps, takes an average of 109 seconds, and requires a peak memory of 3.44 GB. During inference, generating text attributes for a single graph takes an average of 80 seconds. In our experiments, we deploy 8 fine-tuned LLMs in parallel to synthesize 1,000 graphs, completing the process in 2.78 hours. The total GPU memory required for the 8 models is 28 GB. This analysis demonstrates that ProvSyn is a time- and memory-efficient framework for provenance graph synthesis.
6 Related Work
Provenance-based Intrusion Detection System (pids). Provenance-based detection methods can be divided into three main categories [zengy2022shadewatcher]: rule-based, statistics-based, and learning-based techniques. Rule-based methods [hassan2020tactical, milajerdi2019holmes, hossain2020combating] rely on predefined heuristics derived from known attack patterns to identify malicious activities. Statistics-based [hassan2019nodoze, wang2020you, liu2018towards] approaches analyze deviations in provenance graph elements by measuring their statistical anomalies. Learning-based techniques employ deep learning models to capture either normal system behavior [han2020unicorn, wang2022threatrace] or malicious patterns [liu2019log2vec, zengy2022shadewatcher, li2021hierarchical], framing APT detection as either a classification [li2021hierarchical] or anomaly detection task [zengy2022shadewatcher, han2021sigl]. These learning-based methods further branch into sequence-based approaches [liu2019log2vec], which focus on execution workflows, and graph-based [han2020unicorn, wang2022threatrace, zengy2022shadewatcher, han2021sigl, li2021hierarchical] methods, which use graph neural networks to model entity relationships and detect behavioral anomalies.
LLM-driven Synthetic Data Generation. LLM-based synthetic data generation methods can be categorized into three main types: prompt engineering-based approaches [gilardi2023chatgpt, gunasekar2023textbooks, sudalairaj2024lab, liu2024best, yoo2021gpt3mix, li2023textbooks, ye2022progen], which utilize task descriptions, condition-value prompts, and in-context examples to steer generation; multi-step generation-based approaches [wei2022chain, gao2023retrieval, ding2023enhancing, wang2022self, honovich2022unnatural], which break down complex tasks either at the sample or dataset level; and knowledge-enhanced approaches [pan2024unifying, chen2024mindsearch, ji2021survey], which incorporate external knowledge such as knowledge graphs or web sources to improve the factual quality of outputs. The evaluation of synthetic data includes both direct metrics, which measure faithfulness [lee2022factuality] and diversity [yu2023large], and indirect metrics, which assess performance on downstream tasks through benchmark testing [sun2023principle] and open-ended evaluation using human [he2023annollm] or model-based [xu2023wizardlm] methods.
@@ 7 Discussion and Limitations
Temporal Modeling. A provenance graph is not only a directed heterogeneous graph with textual attributes, but also a temporal graph in which each event is associated with a timestamp indicating its occurrence time in the system. These temporal information has recently been leveraged by downstream intrusion detection systems [rehman2024flash] to better capture the patterns of benign and anomalous behaviors. However, in ProvSyn, we do not synthesize explicit timestamps for events. Instead, we approximate temporal order by relying on the sequence of directed edges. Future work may explore modeling temporal dynamics using temporal graph generation networks [zhang2020tg], or designing strategies that leverage LLMs to generate realistic timestamps, thereby enabling more comprehensive and temporally coherent provenance graph synthesis.
Graph Scale. In ProvSyn, the generated provenance graphs are smaller in scale compared to real-world provenance graphs. To enable data augmentation, we embed multiple synthetic graphs as distinct communities within large real-world provenance graphs. The results show that this approach effectively mitigates data imbalance and improves accuracy of detection models trained on the augmented datasets. Future work can focus on scaling up synthetic provenance graphs by enhancing graph generation models for large-scale synthesis and exploring graph merging algorithms that combine smaller graphs into larger ones while preserving structural fidelity.
@@ 8 Conclusion
This paper proposes ProvSyn, a novel provenance graph generation framework consisting of three compoments: graph structure synthesis via heterogeneous graph generation models, textual attribute synthesis via fine-tuned LLM, and five-dimensional fidelity evaluation. Compared to provenance data synthesized directly by four state-of-the-art LLMs, ProvSyn consistently produces graphs with higher fidelity. When integrated into APT detection training sets, ProvSyn-generated graphs effectively mitigate class imbalance and yield improved accuracy.
Ethical Considerations
To the best of our knowledge, this work does not present any ethical concerns. All experiments were conducted on publicly available datasets that were obtained in compliance with ethical standards and do not contain any sensitive or personally identifiable information. The use of large language model APIs in this research adhered strictly to the official terms of service and usage policies provided by the respective providers.
Open Science
To to facilitate future research in provenance graph synthesis and pids, we release the source code of ProvSyn at https://anonymous.4open.science/r/OpenProvSyn-4D0D/.
Appendix A Entity and Event Distribution
@@
We present the distributions of entity type and event type across six commonly-used public provenance datasets in Figure 4. The distributions exhibit a typical long-tail pattern and reveal significant data imbalance.
Appendix B Post-processing Rules
Table 5: Dataset-specific post-processing schemas
Dataset Node Pair (Src, Dst) Permissible Event Types
DARPA E3/E5 (subject, subject) read, write, open, connect, recvfrom, sendto, clone, sendmsg, recvmsg
(subject, file) write, connect, sendmsg, sendto, clone
(subject, netflow) write, sendto, connect, sendmsg
(file, subject) read, open, recvfrom, execute, recvmsg
(netflow, subject) open, read, recvfrom, recvmsg
OpTC (subject, subject) create, open, terminate
(subject, file) create, delete, modify, rename, write
(file, subject) read
(subject, netflow) message, start
(netflow, subject) message, open, start
For each dataset, we define a dataset-specific post-processing schema that constrains the permissible event types between different pairs of node types, as summarized in Table 5. These schemas are derived from the structural regularities observed in the corresponding ground-truth provenance graphs.
Appendix C Baseline LLM Prompt
We present in Figure 7 the prompt used to instruct LLMs in generating provenance logs on Cadets-E3 dataset. In our experiments, the Examples section contains 20 examples. For presentation brevity, only 5 examples are shown.
Appendix D Hyperparameters Setting
Hidden size is of LSTM for graph structure synthesis. Hidden size is a key parameter for LSTM, as it determines the capacity of the model to capture structural dependencies in sequential node and edge generation. We conducted experiments on the Theia-E3 dataset and evaluated structural fidelity using five MMD-based metrics in the validation set. The results in Figure 5 (a) show that a hidden size of 64 yields significantly lower fidelity compared to the other settings. For hidden sizes of 128, 256, and 512, the MMD values remain at a comparable level, with 256 performing slightly better than 128 and 512. Considering both performance and computational cost, we select 256 as the hidden size for ProvSyn.
Inference temperature is a key parameter in LLM-based inference, controlling the randomness of generated text. Higher temperatures yield more diverse outputs, while lower temperatures produce more deterministic results. We conducted experiments on the Cadets-E3 dataset using temperature values of 0.1, 0.5, 1.0, and 1.5, with all other inference parameters held constant. Evaluation focused on the quality of text generated for process and file node attributes. As shown in Figure 5 (b) and (c), the best performance is achieved when the temperature is set to 1.5. Based on these findings, we selected 1.5 as the final inference temperature for ProvSyn.
Appendix E Ablation Study
Alternative models to GraphGen. For heterogeneous graph structure generation, we adopt the design of GraphGen, which encodes graphs as DFS code sequences and employs an LSTM for training and inference. To validate the effectiveness of this design, we compare it with alternative models, including the widely used GraphRNN [you2018graphrnn]. In our experiments, we evaluate the graphs generated by GraphRNN and GraphGen using four MMD metrics. As shown in Figure 6(a), the graphs generated by GraphGen consistently achieve lower MMD scores across all MMD metrics, indicating better fidelity compared to those generated by GraphRNN. These results demonstrate the effectiveness of using GraphGen for heterogeneous graph generation in ProvSyn.
Masking strategies. For text attribute generation, we employ two masking strategies to construct the training set, corresponding to two different input sequences during inference. To evaluate the effectiveness of the masking design, we perform ablation experiments by disabling one of the strategies and using only the other. Model performance is then evaluated using textual and semantic metrics. The experiments are conducted on the Clearscope-E5 dataset, with the number of training samples fixed at 20,000 for both the single-strategy and dual-strategy settings. In the results shown in Figure 6(b), we observe that using only the Full Masking strategy yields better performance than using only the Part Masking strategy. This may be because the Part Masking strategy provides partial node names within the sequence, offering more cues for the LLM to leverage. As a result, a model trained solely with Part Masking performs poorly when tested on Full Masking sequences, which contain fewer clues. In contrast, using both strategies together achieves better performance than either alone, as it exposes the model to a wider range of sequence types that may appear during inference. These results demonstrate the effectiveness of our proposed approach that combines both masking strategies.
LLM parameter size. In our experiments on text attribute generation, we employ the Llama-3.2-3B-Instruct model as the base model for training. We also compare its performance against models of different sizes, including Llama-3.2-1B-Instruct and Llama-3.1-8B-Instruct. All models are trained on the same training datasets and evaluated using textual and semantic metrics. The results, presented in Figure 6(c), demonstrate that the 3B model achieves the highest scores on the Distinct-1 and semantic accuracy. In contrast, the 8B model shows marginally better performance on the GLEU. Given that the advantage of the 8B model is not substantial and considering computational resource constraints, we select the 3B model as the base for synthesizing text attributes. In practice, ProvSyn introduces a general training strategy to leverage LLMs for synthesizing text attributes, where the choice of the base model involves balancing the complexity of the dataset, the application requirements and the limitations of computational resources.
Figure 4: Entity Type and Event Type Distribution in Provenance Dataset.
Figure 5: Hidden size setting and inference temperature setting.